OpenStudy (anonymous):

How do I find the derivative of y=4x-3x^2 at point (2,-4)?

OpenStudy (amistre64):

your going to have to use the rather basic power rule for derivatives

OpenStudy (amistre64):

\[\frac d{dx}(x^n)=n~x^{(n-1)}\]

OpenStudy (anonymous):

How do I solve this by applying the lim/h->0 (f(a+h)-f(a))/h formula?

OpenStudy (amistre64):

with lamenting ...

OpenStudy (anonymous):

I've attempted this problem a lot, and can't figure out how to use the formula to find the an of -8x+12

OpenStudy (anonymous):

*answer

OpenStudy (amistre64):

just replace x with a, then replace x with (a+h) and the algebra tends to take care of itself

OpenStudy (anonymous):

and a in this case will be 2 correct?

OpenStudy (amistre64):

a is just a generic filler

OpenStudy (amistre64):

\[\frac{4(a+h)-3(a+h)^2)-(4a-3a^2)}{h}\]

OpenStudy (amistre64):

in the end, you would fill in the remaining as with 2 yes

OpenStudy (amistre64):

of course I see no reason why you couldnt fill in a at teh start ... just habit i spose :)

OpenStudy (kainui):

@Icecicales do you know what this formula means? \[\lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h }\] It is the same thing as this formula you had in algebra to find the slope, but "calculus" \[\frac{ rise }{ run }=\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\]

OpenStudy (anonymous):

OH,ok, let me try that out

OpenStudy (anonymous):

I've never been been good with math and I have to have everything broken down in the smallest terms to understand it.

OpenStudy (kainui):

I swapped my variables around, but you actually end up with the same answer since -1/-1=1/1 anyways but don't let that confuse you... lol

OpenStudy (amistre64):

or rather, the limit as the distance in the run approaches zero

OpenStudy (anonymous):

Ok, I just plugged in everything a that amistre64 sugggested, but got -3h-20

OpenStudy (amistre64):

anyways: \[\frac{(4(a+h)-3(a+h)^2)-(4a-3a^2)}{h}\] \[\frac{4a+4h-3a^2-6ah-3h^2-4a+3a^2}{h}\] \[\frac{4h-6ah-3h^2}{h}\] \[\frac{h(4-6a-3h)}{h}\] \[4-6a-3h:~h=0\] \[4-6a:~a=2;-8\]

OpenStudy (amistre64):

in essense, the algebra produces and equivalent function to evaluate without the offending zero denominator

OpenStudy (anonymous):

Thank you dude!!!!!!!

OpenStudy (amistre64):

youre welcome, I hope you found the spot that you were missing :)

OpenStudy (anonymous):

Thank you both